# Towards Better Open-Ended Text Generation: A Multicriteria Evaluation Framework

### v1

@@ Abstract
Open-ended text generation has become a prominent task in natural language processing due to the rise of powerful (large) language models. However, evaluating the quality of these models and the employed decoding strategies remains challenging because of trade-offs among widely used metrics such as coherence, diversity, and perplexity. Decoding methods often excel in some metrics while underperforming in others, complicating the establishment of a clear ranking. In this paper, we present novel ranking strategies within this multicriteria framework. Specifically, we employ benchmarking approaches based on partial orderings and present a new summary metric designed to balance existing automatic indicators, providing a more holistic evaluation of text generation quality. Furthermore, we discuss the alignment of these approaches with human judgments. Our experiments demonstrate that the proposed methods offer a robust way to compare decoding strategies, exhibit similarities with human preferences, and serve as valuable tools in guiding model selection for open-ended text generation tasks. Finally, we suggest future directions for improving evaluation methodologies in text generation. Our codebase, datasets, and models are publicly available.
@@ 1 Introduction
Figure 1: Ranking models and decoding strategies, i.e. decoding methods within a multicriteria framework using partial orders and summarizing metrics.
@@
Decoding methods frequently exhibit trade-offs, excelling in some metrics while underperforming in others, complicating the establishment of a consistent ranking among them. Metrics such as MAUVE (Pillutla et al., 2021), coherence (Su et al., 2022), diversity, and generation perplexity (Jelinek et al., 2005) are commonly used to assess the quality of generated text and the generation confidence. While MAUVE measures the alignment between generated and human-like text distributions, coherence evaluates logical flow, diversity ensures varied outputs, and perplexity assesses the predictability of text. While each metric provides valuable insights, they offer incomplete perspectives on the overall generation quality.
Human evaluation, though often considered the gold standard, is resource-intensive and subjective, making it impractical for large-scale assessments of generated texts. The challenge lies in balancing these metrics to create a comprehensive evaluation framework that can reliably compare different models and decoding strategies – collectively referred to as decoding methods in this work.
@@
More LLMs and decoding strategies imply that the number of such decoding methods grows strongly. This development renders reliable benchmarking methods for comparing them even more important. In practice, two prototypical examples of benchmarking scenarios can be distinguished. Firstly, consider a practitioner using open-ended text generation for a very specific task, e.g., a customer support chatbot. This practitioner might primarily be interested in simply knowing which decoding method works best for this particular scenario and may additionally be keen to know which method comes second, third, etc. This motivation renders metric information about the methods’ performances a means to an end. Secondly, think of a researcher who is interested in designing new decoding methods (i.e., either novel models or novel decoding strategies or both). For such a researcher, it is of utmost importance to know how much better one method is compared to another one, i.e., have metric information. Knowing about how existing decoding methods perform on different tasks might help them to derive new methods.
In reality, there are of course many more scenarios in between these two extreme cases. In this work, we propose multiple benchmarking methods along this spectrum. What unites them is their ability to facilitate comparisons across multiple criteria. Specifically, we present approaches based on partial orderings and pairwise comparisons to handle the multicriteria nature of the evaluation and introduce a new summary metric that captures the balance among existing automatic indicators. These approaches simplify the comparison process, enable more robust decoding method selection for open-ended text generation tasks, and effectively balance the trade-offs between competing metrics.
Contributions
@@
  3. 3.
  4. 4.
    We critically present and discuss benchmarking methods for evaluating models and decoding strategies, highlighting their strengths and weaknesses to guide robust decoding method selection.
  5. 5.
    We make our codebase and datasets publicly accessible at this link.
@@ 2 Related Work
Benchmarks are crucial to any kind of applied machine learning research Zhang and Hardt (2024a); Shirali et al. (2023); Ott et al. (2022); Zhang et al. (2020); Thiyagalingam et al. (2022); Roelofs et al. (2019); Vanschoren et al. (2014). On the one hand, they help practitioners make informed choices among different methods for a given problem. On the other hand, newly proposed methods typically need to demonstrate their benefits through such benchmark studies, see for instance Meyer et al. (2003); Hothorn et al. (2005); Eugster et al. (2012); Mersmann et al. (2015). In the last two decades or so, detecting statistically significant differences between methods has attracted increasing attention, recognizing the fact that problems (like datasets) in a benchmark suite are only samples of a population that needs to be defined, see Demšar (2006); García and Herrera (2008); García et al. (2010); Benavoli et al. (2017) for some pioneering works.
@@
Open-ended text generation is no exception. Several metrics to evaluate the quality of decoding strategies have been proposed and discussed in recent years Alihosseini et al. (2019); Pillutla et al. (2021); Celikyilmaz et al. (2021); Su and Xu (2022); Su et al. (2022); Gao et al. (2022); Becker et al. (2024); Garces-Arias et al. (2024a). Diversity, MAUVE, coherence, and generation perplexity are among the most popular metrics. Diversity measures lexical variation using $n$-gram repetition rates, with higher scores indicating less repetition. MAUVE is a distribution similarity metric between generated and reference text. Coherence is defined as the averaged log-likelihood of the generated text conditioned on the prompt. Finally, generation perplexity Jelinek et al. (2005) measures the predictability of the generated text under the language model; lower perplexity indicates that the text is more likely according to the model’s own probability distribution.
This multitude of quality metrics naturally gives rise to the question of how to aggregate them, i.e., how to compare decoding methods holistically, taking into account multiple dimensions of text quality. It is self-evident that focusing on only one metric has obvious shortcomings. For instance, exclusively optimizing coherence will favor decoding methods with only moderate to low penalties for already selected tokens. This can lead to repetitive and uncreative text generation; a problem often referred to as degeneration Lee et al. (2022); Holtzman et al. (2019). On the other hand, focusing solely on diversity will eventually result in incoherent text only slightly – if at all – related to the prompt. In this work, we address exactly the problem of how to benchmark decoding strategies with respect to multiple and potentially contradicting criteria. In particular, we offer a fresh perspective on this problem, adopting recent developments in the theory of depth functions, see section 4. Moreover, we propose a novel summary metric to balance multiple metrics, see section 5.
@@ 3.2 Evaluation Metrics
Building upon Su and Collier (2023), we employ various automatic metrics to assess the quality of the generated texts: Diversity, Coherence, and Perplexity. Additional metrics like the MAUVE score measure the distributional similarity between generated texts and reference texts. However, since MAUVE relies solely on aggregated data and requires reference texts, we exclude it from our evaluations in this study.

### v2

@@ Abstract
Open-ended text generation has become a prominent task in natural language processing due to the rise of powerful (large) language models. However, evaluating the quality of these models and the employed decoding strategies remains challenging because of trade-offs among widely used metrics such as coherence, diversity, and perplexity. Decoding methods often excel in some metrics while underperforming in others, complicating the establishment of a clear ranking. In this paper, we present novel ranking strategies within this multicriteria framework. Specifically, we employ benchmarking approaches based on partial orderings and present a new summary metric designed to balance existing automatic indicators, providing a more holistic evaluation of text generation quality. Our experiments demonstrate that the proposed methods offer a robust way to compare decoding strategies, and serve as valuable tools in guiding model selection for open-ended text generation tasks. Finally, we suggest future directions for improving evaluation methodologies in text generation. Our codebase, datasets, and models are publicly available. [^1] https://github.com/YecanLee/2BeOETG
@@ 1 Introduction
Figure 1: We present a multicriteria evaluation framework for benchmarking models and decoding strategies, i.e., decoding methods. We distinguish two scenarios for benchmarking (Section 1) and two ranking objectives (Section 4), giving rise to three use-case tailored, distinct methods (sections 4.14.2 and 5).
@@
Decoding methods frequently exhibit trade-offs, excelling in some metrics while underperforming in others, complicating the establishment of a consistent ranking among them. Metrics such as MAUVE (Pillutla et al., 2021), coherence (Su et al., 2022), diversity, and generation perplexity (Jelinek et al., 2005) are commonly used to assess the quality of generated text and the generation confidence. While MAUVE measures the alignment between generated and human-like text distributions, coherence evaluates logical flow, diversity ensures varied outputs, and perplexity assesses the predictability of text. While each metric provides valuable insights, considering them in isolation offers an incomplete perspectives on the overall generation quality. Human evaluation, though often considered the gold standard, is resource-intensive and subjective, making it impractical for large-scale assessments of generated texts. In this work, we focus on automatic evaluation of generated text. The challenge lies in balancing multiple automatic (and often contradictory) metrics to create a comprehensive evaluation framework that can reliably compare different models and decoding strategies – collectively referred to as decoding methods in this paper. More LLMs and decoding strategies imply that the number of such decoding methods grows strongly. This development renders reliable benchmarking methods for comparing them even more important.
In practice, two prototypical examples of benchmarking scenarios can be distinguished, see Fig. 2.
@@
Scenario 1: Ranking Firstly, consider a practitioner using open-ended text generation for a very specific task, e.g., a customer support chatbot. This practitioner might primarily be interested in simply knowing which decoding method works best for this particular scenario and may additionally be keen to know which method comes second, third, etc. This motivation renders metric information about the methods’ performances a means to an end. Thus, a ordinal ranking of methods will do, see Figure 2 (top).
Scenario 2: Metric Assessment Secondly, think of a researcher who is interested in designing new decoding methods (i.e., either novel models or novel decoding strategies or both). For such a researcher, it is of utmost importance to know how much better one method is compared to another one, i.e., have metric information, see Figure 2 (bottom). Knowing about how existing decoding methods perform on different tasks might help them to derive new methods.
In reality, there are of course many more scenarios in between these two prototypical cases. In this work, we propose multiple benchmarking methods along this spectrum. What unites them is their ability to facilitate multiple criteria. This way, we address the following issues.
Research Gap: When evaluating decoding methods based on multiple criteria (such as metrics or tasks) across several test problems (like different datasets), a method may excel in one area while lagging in another. It is still understood as an open problem of how to aggreate such multicriteria evaluation results in different scenarios, see Figure 1 and 2. Existing approaches comprise the Pareto front – all decoding methods that are undominated across all criteria – or weighted sums of the criteria. While the former wastes a lot of information, see also Section 4, the latter relies arbitrarily on the selected weights. This paper offers alternatives. In particular, we address the following questions.
Research Questions
@@
    RQ1: Can we exploit novel statistical methodologies for partial orders to establish multicriteria rankings that potentially allow for incomparability? (Scenario 1: Ranking)
@@
    RQ2: Can we aggregate multiple automatic evaluation metrics in a meaningful and statistically valid way? (Scenario 2: Metric Assessment)
We address RQ1 in Section 4 and RQ2 in Section 5 by proposing appropriate aggregation methods, see also Figure 1. What is more, we apply all these methods on 1.8 million stories generated by six state-of-the-art models on real world datasets, see Section 3 (setup) and Section 6 (results).
Figure 2: Illustration of ordinal ranking (top) and metric comparison (below) of decoding methods (Scenarios 1 and 2, respectively). Top: Distances are not meaningful, lower three methods are incomparable. Bottom: Distances are meaningful, all methods are comparable.
@@ 2 Related Work
Benchmarks are the backbone of any kind of applied machine learning research Zhang and Hardt (2024a); Shirali et al. (2023); Ott et al. (2022); Zhang et al. (2020); Thiyagalingam et al. (2022); Roelofs et al. (2019); Vanschoren et al. (2014). On the one hand, they help practitioners make informed choices among different methods for a given problem. On the other hand, newly proposed methods typically need to demonstrate their benefits through such benchmark studies, see for instance Meyer et al. (2003); Hothorn et al. (2005); Eugster et al. (2012); Mersmann et al. (2015). In the last two decades or so, detecting statistically significant differences between methods has attracted increasing attention, recognizing the fact that problems (like datasets) in a benchmark suite are only samples of a population that needs to be defined, see Demšar (2006); García and Herrera (2008); García et al. (2010); Benavoli et al. (2017) for some pioneering works.
@@
This shift towards multi-citeria/task benchmarking is particularly prevalent in natural language processing (NLP). Modern generalist models, like LLMs, are evaluated with respect to a multitude of metrics, owing to their breadth and depth of capabilities, see e.g., Wei et al. (2024); Liu et al. (2025). As we can only improve upon what we measure, multi-criteria benchmarking has moved to the forefront of scientific inquiry. By assessing models across diverse tasks – ranging from reasoning and comprehension to creativity and ethical considerations – researchers gain a more holistic understanding of their strengths and limitations Chiang et al. (2024). Such comprehensive evaluation frameworks not only drive advancements in model performance but also ensure alignment with real-world applications and ethical standards Liu et al. (2023); Ji et al. (2023); Terry et al. (2023); Rodemann et al. (2025). As a result, multi-criteria benchmarking has become an essential pillar in the evolution of AI, guiding both theoretical progress and practical deployment.
Decoding methods for open-ended text generation are no exception. Several metrics to evaluate the quality of decoding strategies have been proposed and discussed in recent years Alihosseini et al. (2019); Pillutla et al. (2021); Celikyilmaz et al. (2021); Su and Xu (2022); Su et al. (2022); Gao et al. (2022); Becker et al. (2024); Garces-Arias et al. (2024a, 2025). Diversity, MAUVE, coherence, and generation perplexity are among the most popular metrics. Diversity measures lexical variation using $n$-gram repetition rates, with higher scores indicating less repetition. MAUVE is a distribution similarity metric between generated and reference text. Coherence is defined as the averaged log-likelihood of the generated text conditioned on the prompt. Finally, generation perplexity Jelinek et al. (2005) measures the predictability of the generated text under the language model; lower perplexity indicates that the text is more likely according to the model’s own probability distribution.
This multitude of quality metrics naturally gives rise to the question of how to aggregate them, i.e., how to compare decoding methods holistically, taking into account multiple dimensions of text quality. It is self-evident that focusing on only one metric has obvious shortcomings. For instance, exclusively optimizing coherence will favor decoding methods with only moderate to low penalties for already selected tokens. This can lead to repetitive and uncreative text generation; a problem often referred to as degeneration Lee et al. (2022); Holtzman et al. (2019). On the other hand, focusing solely on diversity will eventually result in incoherent text only slightly – if at all – related to the prompt.
In this work, we address exactly the problem of how to benchmark decoding strategies with respect to multiple and potentially contradicting criteria. In particular, we offer a fresh perspective on this problem, adopting recent developments in the theory of depth functions and order theory, see Section 4. Moreover, we propose a novel summary metric to balance multiple metrics, see Section 5.
@@ 3.2 Evaluation Metrics
Building upon Su and Collier (2023), we employ various automatic metrics to assess the quality of the generated texts: Diversity, Coherence, and Generation Perplexity (Jelinek et al., 2005; Holtzman et al., 2019). Additional metrics like the MAUVE score measure the distributional similarity between generated texts and reference texts. However, since MAUVE relies solely on aggregated data and requires reference texts, we exclude it from our evaluations in this study. A detailed overview of their definitions is provided in Section A.1.
3.3 Datasets
Furthermore, we utilize the dataset provided by Garces-Arias et al. (2024a), which includes over 2.2 million generated continuations (with a maximum length of 256 tokens) for each prompt, along with aggregated metrics such as coherence, diversity, and MAUVE. We extend this dataset by computing sentence-level evaluations and incorporating generation perplexity, as described above in Section 3.2.
3.4 Decoding Strategies and Hyperparameters
4 Benchmarking utilizing Order Structure
Characteristic Extended Bradley-Terry Model Union-Free Generic Depth Q*Text
Metric information considered Order only Order only Order and metric value
Methodology basis Pairwise comparison Partial orders Mean values
Output Returns a worth parameter and a total order Returns a partial order Returns mean values and a total order
Results (based on WikiText-103) Mistral 3 CS ((’0.4’, ’10’)) has the highest worth parameter, while GPT2-XL CS ((’1.0’, ’20’)) has the lowest. The top five models in the Extended Bradley-Terry Model are incomparable, despite the model suggesting a total order. Falcon 2 CS ((’0.8’, ’1’)) has the highest mean and Mistral 3 CS ((’0.2’, ’1’)) the lowest.
Table 2: Comparison of the three methods: extended Bradley-Terry, ufg-depth and Q*Text, see also Figure 1.
In this section, we propose benchmarking of decoding strategies along the Bradley-Terry model Bradley and Terry (1952b) and the union-free-generic (ufg) depth Blocher et al. (2024); Blocher and Schollmeyer (2024). Both approaches deliver rankings of decoding strategies rather than metric values for each strategy. We motivate this from a very practical perspective, see Section 1: In many cases, the metric information can be considered redundant. That is, ranking the decoding strategies is the overall aim of benchmarking, not assigning metric values. After all, a decoding method can either be deployed or not by practitioners, rendering the metric information not of direct practical interest.
4.1 Extended Bradley-Terry Model
The extended Bradley-Terry model is based on pairwise comparisons Bradley and Terry (1952a); Davidson (1970). To obtain this pairwise comparison of the methods, we compare the methods pairwise on each single generation. For simplicity, let us look at WikiText-103 with 1314 generations. For each generation $G_{i,j}$ (from model $i$ with decoding strategy $j$, see Section 1), we can evaluate the methods based on coherence, perplexity, and diversity. Now we consider each pair of methods and each generation, e.g. Mistral 3 CS ((’0.2’, ’1’)) and Mistral 3 CS ((’0.8’, ’1’)) with generation id 1. Then we state that method 1 dominates method 2 if and only if there exists one metric that strictly prefers method 1 over method 2 and the other two metrics do not disagree. Based on generation id 1, we obtain that Mistral 3 CS ((’0.2’, ’1’)) is preferred over Mistral 3 CS ((’0.8’, ’1’)) since coherence states that method Mistral 3 CS ((’0.2’, ’1’)) is preferred over method Mistral 3 CS ((’0.8’, ’1’)), and the other two metrics, here diversity and perplexity, do not contradict this statement. In this way, for WikiText-103 we get 1314 comparison statements for each pair of two methods. Note that when two methods have exactly the same metric values, they are considered indifferent. This differs from incomparability; for details, see Rodemann and Blocher (2024). However, with slight abuse, we do not distinguish these two cases here.
4.2 Union-Free Generic Depth
To account for exactly this complexity, we build on Rodemann and Blocher (2024); Blocher et al. (2024) and deliberately allow for incomparability. Mathematically, this means that we aim for partial rather than total orders. Let us look again at a single generation $G_{i,j}$. For the extended Bradley-Terry model, we only considered the pairwise comparisons. However, all the pairwise comparisons resulting from a generation $G_{i,j}$ together define a partial order that describes the preference structure of the methods based on that generation $G_{i,j}$. This gives us 1314 partial orders for the WikiText-103 data. For example, in the case where we compare four methods, the two partial orders in Figure 3 correspond to two observation.
5 Benchmarking by Collapsing into a Single Metric
As previously discussed, metrics such as coherence and diversity, when optimized excessively, can lead to unintended outcomes. High coherence scores often result in repetitive or degenerate text, as the model overemphasizes continuity at the expense of introducing new information. Conversely, prioritizing high diversity can produce erratic sequences that lose relevance to the initial prompt, undermining the overall quality of the generated text. Q*Text consists of the harmonic mean of normalized and penalized coherence, diversity and generation perplexity. Its construction involves three key steps: normalization to a common scale of 0 to 100, where 0 represents the worst performance and 100 represents the best, penalization of extreme values, and aggregation using the harmonic mean. A technical description is detailed in Section A.2.
6 Results
Extended Bradley-Terry Model
The extended Bradley-Terry model returns so-called "worth" parameters which indicate the probability that this decoding method is preferred over the other in a pairwise comparison. When all datasets are considered at once, the method that dominates all other methods according to the extended Bradley-Terry model is Mistral 3 CS ((’0.6’, ’15’)). The second best method is Mistral 3 CS ((’0.4’, ’5’)). The worst method is GPT2-XL CS ((’1.0’, ’20’)). A snippet of the results, including the case when restricting the analysis to only one dataset, can be found in Table 3.
Method Estimated worth parameter
Mistral 3 CS ((’0.6’, ’15’)) 0.046 940 673 747 754 145 528 343 627 802 314 586 006 104 946 136 470.046940673747754145528343627802314586006104946136470.046\,940\,673\,747\,754\,145\,528\,343\,627\,802\,314\,586\,006\,104\,946\,136\,470.046 940 673 747 754 145 528 343 627 802 314 586 006 104 946 136 47
Mistral 3 CS ((’0.4’, ’3’)) 0.037 445 510 668 860 391 978 451 929 162 474 698 387 086 391 448 970.037445510668860391978451929162474698387086391448970.037\,445\,510\,668\,860\,391\,978\,451\,929\,162\,474\,698\,387\,086\,391\,448\,970.037 445 510 668 860 391 978 451 929 162 474 698 387 086 391 448 97
Mistral 3 CS ((’0.8’, ’3’)) 0.034 596 596 197 846 059 272 951 180 219 024 536 199 867 725 372 310.034596596197846059272951180219024536199867725372310.034\,596\,596\,197\,846\,059\,272\,951\,180\,219\,024\,536\,199\,867\,725\,372\,310.034 596 596 197 846 059 272 951 180 219 024 536 199 867 725 372 31
Mistral 3 CS ((’0.4’, ’20’)) 0.029 517 106 894 895 648 455 667 824 805 459 531 489 759 683 609 010.029517106894895648455667824805459531489759683609010.029\,517\,106\,894\,895\,648\,455\,667\,824\,805\,459\,531\,489\,759\,683\,609\,010.029 517 106 894 895 648 455 667 824 805 459 531 489 759 683 609 01
Table 3: Estimated worth parameter of the extended Bradley-Terry model based on WikiText-103 dataset and the metric coherence, diversity and perplexity.
Union-Free Generic Depth
Therefore, in the next step, we consider the union-free generic depth approach, which allows for two methods to be incomparable. Furthermore, the ufg-depth considers the entire set of pairwise comparisons for a generation as one observation and does not assume an independence structure between them. Due to the high computational complexity, see Appendix A.7, we restrict our analysis to the WikiText-103 dataset and compare only the four methods that appear to be the best according to the extended Bradley-Terry model, see Appendix A.4: Mistral 3 CS ((’0.6’, ’15’)), Mistral 3 CS ((’0.4’, ’3’)), Mistral 3 CS ((’0.8’, ’3’)) and Mistral 3 CS ((’0.4’, ’20’)).
The highest ufg-depth with a value of 0.977 (thus the one that has the structure most supported by the observation), is the one that shows no dominance structure among the four methods, i.e. the one that concludes that all methods are incomparable to each other, see Figure 3 (top). Roughly speaking, our method reveals that the four decoding methods considered here are incomparable. More formally put, we identify a trivial ranking with no dominance structure as the “central” (in the sense of being the “median”) of the dataset comprising the benchmarking results. This means that such a ranking has most support by the benchmarking results. Our method further finds an “outlier”, i.e., a ranking of methods that has least support by the benchmarking results. In the example at hand, this outlier is a partial ranking that ranks Mistral 3 CS ((’0.4’, ’3’)) higher than Mistral 3 CS ((’0.4’, ’20’)), see Figure 3 (bottom). This means that given the benchmarking results, such a ranking of methods is “least central” or “atypical” and therefore based on the benchmarking results the least supporter structure.
Q*Text
With regards to Q*Text, we observe that balanced deterministic strategies such as contrastive search with moderate penalties ($\alpha$ values of 0.4 or 0.6) and moderate values of $k$ (5, 10, or 15) are preferred, as illustrated in Tables 18 and 16. Combinations that counterbalance each other also yield favorable results – for instance, low $\alpha$ values (0.2) with high $k$ values (20 or 50), or high $\alpha$ values (0.8 or 1.0) with moderate $k$ values (3 or 5). Both Q*Text and the extended Bradley-Terry model indicate a general preference for larger architectures; however, smaller models like GPT2-XL can outperform more modern architectures when paired with balanced decoding strategies, see Section 12.
To evaluate the level of agreement between the total ordering from the extended Bradley-Terry and that provided by Q*Text, we analyze and visualize the decoding methods with the highest agreement and those with the largest rank deviations (see Appendix A.6, Figures 4 and 5). The results suggest discrepancies, particularly for less diverse and less coherent generations. Nevertheless, there is good agreement when ranking decoding methods with moderate hyperparameters, as previously discussed.
Despite operating on different underlying mechanisms, the preferences from the extended Bradley-Terry model exhibit noteworthy similarities with those from Q*Text concerning the aforementioned balanced strategies. One notable difference is that the extended Bradley-Terry model does not account for significant drops in diversity, as seen with beam search, which reduces the penalization of this decoding strategy. On the other hand, decoding methods such as GPT2-XL with contrastive search parameters ($\alpha$ of 1.0 and $k$ of 20) are strongly penalized by both approaches due to their incoherence and low generation confidence, see Tables 13, 15 and 19.
Finally, we naturally ask how these alternative methods, evaluated within a multicriteria framework, align with human preferences. To address this, we considered the rankings provided by human evaluators in Section 5 of Garces-Arias et al. (2024a), which present similar findings. Specifically, human evaluators favored balanced contrastive search strategies, e.g., (0.4, 10), (0.4, 20), (0.4, 5), (0.6, 10), and (0.6, 5), top-$k$ sampling with $k=10$ or 50, and nucleus (top-$p$) sampling with $p=0.9$. In contrast, unbalanced contrastive search approaches, such as (1.0, 50) and (0.8, 50), were consistently evaluated as low quality due to their low coherence. Beam search strategies with moderate beam widths, e.g. less than or equal to 20, were also often judged as low quality due to their low diversity. See Tables 10, 11 and 18 and the work by Garces-Arias et al. (2024a) for more details. Furthermore, Table 20 presents examples of generations from various decoding methods, highlighting those that received positive and negative evaluations from Bradley-Terry and Q*Text.
7 Conclusion
By applying these methodologies in extensive experiments across diverse datasets and models, we have shown that our proposed approach not only enhances the interpretability of text generation evaluations but also provides a practical, data-driven framework for benchmarking algorithms in this domain. We hope that these insights will contribute to a more rigorous evaluation of text generation models and inspire further research in the field of multicriteria benchmarking strategies. Summing up, this work makes the following contributions.
Contributions
  1. 1.
  2. 2.
    In order to identify which benchmarking technique is best for which practical use case, we discover two main benchmarking scenarios – a distinction that might be of indepedent interest for future research.
  3. 3.
  4. 4.
  5. 5.
    We critically discuss benchmarking methods for evaluating models and decoding strategies with clear recommendations for practitioners (Section 8).
  6. 6.
    We make our efficiently (see Appendix A.7) implemented benchmarking methods and datasets publicly accessible at this link.
8 Key Takeaways and Practical Recommendations
[…]
Specifically, we identify two crucial questions to be answered beforehand:
  1. 1.
    Is it sufficient to rank methods, or is metric information about the methods’ performances required? (Scenario 1 and 2 in Section 1)
  2. 2.
    Does the use case require a total or partial ordering method, i.e., should the evaluation allow for incomparability among some methods, or should it enforce comparability of all methods? (Section 4)
In case metric information is required and comparability of all methods should be enforced, we recommend our novel aggregation metric Q*Text, see Section 5. If the metric information is not the overall aim, but comparability should still be enforced, we recommend using the Bradley-Terry model, see Section 4.1. Eventually, if a ranking is required that allows for incomparability, we recommend deploying ufg-depth; see Section 4.2.
Limitations
Ethics Statement
Acknowledgments
Currently not added due to anonymization purposes.
Appendix A Appendix
A.1 Automatic metrics